# LawChain: Modeling Legal Reasoning Chains for Chinese Tort Case Analysis

### v1

@@ Introduction
We evaluate eight state-of-the-art LLMs on the LawChaineval benchmark, including GPT-4o (Hurst et al. 2024), o3-mini (OpenAI 2025), Claude-Sonnet-4 (Anthropic 2025), DeepSeek-V3 (Liu et al. 2024), DeepSeek-R1 (Guo et al. 2025), Qwen-3-8B (Yang et al. 2025), InternLM-3-8B (InternLM 2025), Llama-3.1-8B (Grattafiori et al. 2024), ensuring broad coverage of model architecture families. We experiment with two inference-only settings for all models, zero-shot and legal prompting, and three training-based settings for open-source models [^2] Although DeepSeek-R1 and DeepSeek-V3 are open-source models, we experimented via API calls due to computational resource considerations, and did not perform fine-tuning for them.: SFT with syllogism data in Deng et al. (2023), SFT with LawChain-style reasoning data and DPO with LawChain-style preference data.
@@
Legal prompting (denoted as PromptLC). This prompting approach incorporates structured LawChain-style reasoning chains, guiding models to produce responses that reflect multi-step legal reasoning. This approach is applicable across LLMs without the need for model training.
SFTSyll. This setting involves supervised fine-tuning using syllogism-style legal reasoning data provided in Deng et al. (2023). In this approach, training instances are structured into the syllogistic form: a major premise representing applicable legal rules, a minor premise representing case facts, and a conclusion stating the legal outcome.
SFTLC. Using the curated LawChaintrain data, we perform supervised fine-tuning on three open-source models, including Qwen-3-8B, InternLM-3-8B and Llama-3.1-8B.
DPOLC. In the DPO setting, we use the LawChaintrain data as the chosen samples. We further generate corresponding rejected samples by collecting answers generated by zero-shot versions of three models (Qwen-3-8B, InternLM-3-8B and Llama-3.1-8B) on LawChaintrain. The resulting (chosen, rejected) sample pairs are used for DPO training.
@@
SFT Experiments. Models are fine-tuned with a batch size of 16 and a learning rate of 1e-4 with AdamW optimizer (Loshchilov and Hutter 2017). Low-Rank Adaptation (Hu et al. 2022, LoRA) is used for parameter-efficient adaptation. We use 10% of the data held out for validation; the optimal checkpoint is selected based on validation set performance. The three models, Qwen-3-8B, InternLM-3-8B and Llama-3.1-8B, are trained for 120, 160 and 80 steps, respectively. All SFT experiments are carried out using LlamaFactory (Zheng et al. 2024) on 8 NVIDIA A800 GPUs.
@@
Evaluation Results on LawChaineval
We evaluate a broad range of state-of-the-art LLMs on LawChaineval. The evaluation results are reported in Table 1. In the zero-shot setting, most models exhibit limited capability in reasoning over the multiple dimensions of tort legal analysis. The best-performing model without any task-specific prompt or fine-tuning is Qwen-3-8B, achieving an overall score of 48.79. Notably, as one of the most widely used LLMs, GPT-4o achieves only 41.17 overall, underscoring the challenge posed by the tort legal reasoning task.
We observe that the LawChain-style legal prompting approach (PromptLC) leads to consistent performance improvements across all models. The largest gain is observed for DeepSeek-R1, achieving a 34.1% increase over its zero-shot baseline and yielding the highest overall performance among all tested approaches. These results highlight the effectiveness of explicit, structured legal prompting in eliciting enhanced legal reasoning abilities from current models.
We further examine the impact of fine-tuning with reasoning-enhanced data, comparing models trained on syllogistic reasoning data (SFTSyll), LawChain-style reasoning data (SFTLC) and LawChain-style preference data (DPOLC). We find that SFTSyll yields reduced performance compared to zero-shot baselines for all models. This suggests that generic syllogistic reasoning is insufficient to capture the domain-specific structure and nuances in tort reasoning. In contrast, both SFTLC and DPOLC exhibit consistent performance improvements across models, with significant gains [^3] In this paper, significance is tested through t-test ($\alpha$=0.01). observed for Qwen-3-8B and Llama-3.1-8B.
@@
Figure 3: Evaluation results of GPT-4o and DeepSeek-R1 on LawChaineval, categorized by four types of tort liability.
@@
To evaluate the generalizability of our proposed reasoning-incorporated approaches to other tasks, we conduct experiments on two related legal AI tasks: Legal Named-Entity Recognition and Criminal Damages Calculation (Fei et al. 2024). Although both tasks are situated within the context of criminal cases, they share certain similarities with the legal element identification and the compensation calculation components of our LawChain framework. We compare the performance of reasoning-enhanced models trained using SFT and DPO to their zero-shot counterparts. As shown in Table 2, both the SFT-based and DPO-based approaches consistently achieve comparable or better performance on these tasks, suggesting that explicitly integrating legal reasoning chains into model training can enhance model performance across a broader spectrum of legal tasks.
@@
To analyze model performance across different tort categories, we classify test instances in LawChaineval according to their underlying dispute types. We focus on the four most prevalent types of tort liability in real-world legal practice (China Judgments Online 2013): liability for motor vehicle traffic accident, liability for injury to providers of labor services, liability for product and liability for medical damage, and examine model performance within each category.
@@ Conclusion
In this work, we propose LawChain, a structured reasoning framework that explicitly models legal reasoning process for tort-related cases within the context of Chinese civil law. We introduce a benchmark to evaluate LLMs’ ability to identify key reasoning steps in tort case analysis, and find that existing models exhibit notable limitations on this task. By incorporating LawChain-style reasoning through legal prompting, supervised fine-tuning and direct preference optimization, we observe consistent performance improvements on tort reasoning. These reasoning-enhanced models also generalize well to other legal AI tasks. These results highlight the importance of explicitly modeling legal reasoning structures to enhance LLMs’ ability for authentic, domain-sensitive legal analysis.
Appendix A Legal Element Extraction
To construct the evaluation dataset LawChaineval and the training dataset LawChaintrain, a crucial preparatory step is to extract key legal elements from original court decisions. For elements commonly expressed in structured formats in court decisions, such as parties (plaintiffs and defendants), references of legal provisions and final judgments, we employ rule-based extraction techniques using keyword matching and regular expression-based heuristics. For elements that require document-wise summarization, such as facts, liability reasoning and damages calculation, we utilize DeepSeek-V3 to synthesize relevant information from the court decisions.
When using DeepSeek-V3 for legal element extraction, the temperature is set to 0.7 and the maximum token size is set to 4096 tokens. As an example, the prompt (English translation) used for extracting fact descriptions of the case from the original court decision is presented in Table 3.
Appendix B Examples of LawChain-Style Training Data
In Table 4, we present the construction rules and illustrative examples of LawChain-style reasoning texts (shown in English translation). The reasoning texts corresponding to different elements of a case are concatenated to form a comprehensive reasoning chain, which is further encapsulated within a <<think>> and <</think>> tag pair to serve as reasoning-enhanced training data.
Appendix C Further Analysis of Tort Liability Types
In Figure 4, we present evaluation results for all zero-shot models across four representative categories of tort cases, including liability for motor vehicle traffic accident, liability for injury to providers of labor services, liability for product and liability for medical damage. These categories correspond to the most prevalent and consequential forms of tort liability encountered in real-world legal practice, collectively covering a broad spectrum of harms that arise in civil litigation.
Across all evaluated models, we observe clear differences in model performance between these categories. Notably, disputes involving injury to providers of labor services consistently emerge as the most challenging case type, exhibiting the lowest average scores regardless of model architecture. This pattern likely reflects the inherent factual and legal complexity of labor-related torts, which often require nuanced reasoning about employment relationships, duty of care, and allocation of liability among multiple parties. By contrast, cases involving traffic accidents, while still non-trivial, tend to feature more standardized fact patterns and established legal doctrines, contributing to relatively higher model performance in this category.
These findings highlight the importance of case-type stratification when evaluating LLM legal reasoning capabilities, and underscore the need for further research into modeling the intricate reasoning required in complex, labor-related tort scenarios.
Element Prompt
Facts You are a legal assistant. Based on the court decision, extract the case facts according to the following requirements:
1. Extract only objective facts confirmed by the court (time, place, action, result). Only extract content after “The Court Finds” section.
2. If the fact is identified through investigation, appraisal, or court commission, retain that qualifier. For example: “This accident was determined by the Linqu County Public Security Bureau Traffic Police Brigade.”
3. The names of the parties (e.g., “Plaintiff XXX” / “Defendant XXX”) must be fully preserved.
4. Fully preserve details about amounts, quantities, units and other numeric information.
5. Number each extracted fact, then output as a list of fact descriptions.
Table 3: Prompt used for extracting fact descriptions from original court decisions (shown in English translation).
Construction Rule Reasoning Text
(Facts, Claims) $\to$ Parties Now, I will determine the parties to the lawsuit based on the facts and the plaintiffs’ claims. The plaintiffs are Plaintiff A and Plaintiff B, the parents of the deceased. They explicitly requested compensation in their complaint. The defendant is Defendant C, which, by unlawfully piling construction materials at the site of the incident, caused the accident and was thus named by the plaintiffs as the party to be held liable.
(Facts, Claims) $\to$ Dispute Next, I need to analyze and determine the cause of action based on the facts and the plaintiffs’ claims. The plaintiffs allege that the defendant unlawfully placed construction materials on a public road without warning signs, resulting in their son’s fatal traffic accident. Such conduct, namely causing harm to others by placing obstructions on a public roadway, falls within the purview of Article 89 of the Tort Liability Law. Therefore, the cause of action in this case is identified as a dispute over liability for obstruction of passage on a public road.
(Facts, Claims) $\to$ Statutes Subsequently, I must reconstruct and identify the applicable legal provisions based on the facts and the plaintiffs’ claims. Three levels of legal application are involved: First, regarding the safety obligation of the road management party, Article 89 of the Tort Liability Law on obstruction of public roads applies. Second, the issue of apportionment of liability in the context of contributory negligence requires application of Article 26 of the Tort Liability Law. Third, the calculation standards for damages are governed by the relevant judicial interpretations: Article 17 (funeral expenses and death compensation) and Article 22 (transportation expenses) of the Interpretation on Compensation for Personal Injury, and Article 10 of the Interpretation on Compensation for Emotional Distress regarding discretionary factors for consolation money.
(Facts, Claims, Statutes) $\to$ Liability_Ratio Next, based on the facts, the plaintiffs’ claims, and the applicable legal provisions, I need to clarify the proportion and type of liability borne by each party. The investigation revealed that the deceased was guilty of drunk driving, speeding, and failing to take evasive measures, constituting triple negligence. Under Article 26 of the Tort Liability Law, this warrants a reduction of the defendant’s liability. Conversely, the defendant, as the construction company, failed to set up warning signs, violating Article 32 of the Road Traffic Safety Law, and should bear corresponding liability under Article 89 of the Tort Liability Law. Taking into account the degree of fault of both parties, the deceased should bear primary responsibility (70%), and the defendant secondary responsibility (30%).
Facts $\to$ Damages For the awarded compensation items, the specific amounts are calculated as follows: Death compensation is calculated at the urban standard for 20 years: RMB 34,757.2 × 20 years = RMB 695,144; Funeral expenses, based on six months of the average monthly wage of employees in Guangdong Province, total RMB 36,329.5; Transportation expenses are set at RMB 500. The sum of these three items constitutes the total material damages of RMB 731,973.5. The consolation money of RMB 20,000 is calculated separately and not subject to apportionment.
(Liability_Ratio, Damages) $\to$ Apportionment Combining the proportion of liability with the total compensation, the defendant’s liability is determined as follows: For material damages: 731,973.5 × 30% = RMB 219,592.05; For consolation money: full amount of RMB 20,000. After deducting the RMB 38,000 already paid by the defendant, the final amount payable is RMB 201,592.05. The court fee of RMB 4,096 is apportioned according to the degree of responsibility: RMB 2,096 by the plaintiffs, RMB 2,000 by the defendant.
(Claims, Apportionment) $\to$ Judgment The final judgment is as follows: 1. The defendant, Defendant C, shall compensate the plaintiffs RMB 201,592.05. 2. The plaintiffs’ claims for support payments and lost wages are dismissed. 3. The court fee of RMB 4,096 shall be borne RMB 2,096 by the plaintiffs and RMB 2,000 by the defendant.
Table 4: Details of LawChain-style training data construction. The left column presents the construction rubrics used to automatically generate LawChain-style reasoning texts based on extracted legal elements. The right column presents the resulting reasoning texts (shown in English translation) as provided in LawChaintrain.
Figure 4: Evaluation results of zero-shot models on LawChaineval, categorized by tort types.

### v2

@@ Introduction
We evaluate current LLMs on LexChaineval, including GPT-4o (hurst2024gpt_4o), o3-mini (openai2025openai_o3mini), Claude-Sonnet-4 (anthropic2025claude_4), DeepSeek-V3 (liu2024deepseek_v3), DeepSeek-R1 (guo2025deepseek_r1), Qwen-3-8B (yang2025qwen_3), InternLM-3-8B (internlm3), and Llama-3.1-8B (grattafiori2024llama_3.1), ensuring broad coverage of model architecture families. We experiment with two inference-only settings for all models, zero-shot and legal prompting, and three training-based settings for open-source models [^2] Although DeepSeek-R1 and DeepSeek-V3 are open-source models, we experimented via API calls due to computational resource considerations, and did not perform fine-tuning for them.: SFT with syllogism data in deng2023syllogistic, SFT with LexChain-style reasoning data and DPO with LexChain-style preference data.
@@
Legal prompting (denoted as PromptLC). This prompting approach incorporates structured LexChain-style reasoning chains, guiding models to produce responses that reflect multi-step legal reasoning. This approach is applicable across LLMs without the need for model training.
SFTSyll. This setting involves supervised fine-tuning using syllogism-style legal reasoning data provided in deng2023syllogistic. In this approach, training instances are structured into the syllogistic form: a major premise representing applicable legal rules, a minor premise representing case facts, and a conclusion stating the legal outcome.
SFTLC. Using the curated LexChaintrain data, we perform supervised fine-tuning on three open-source models, including Qwen-3-8B, InternLM-3-8B and Llama-3.1-8B.
DPOLC. In the DPO setting, we use the LexChaintrain data as the chosen samples. We further generate corresponding rejected samples by collecting answers generated by zero-shot versions of three models (Qwen-3-8B, InternLM-3-8B and Llama-3.1-8B) on LexChaintrain. The resulting (chosen, rejected) sample pairs are used for DPO training.
@@
SFT Experiments. Models are fine-tuned with a batch size of 16 and a learning rate of 1e-4 with AdamW optimizer (loshchilov2017decoupled_adamw). Low-Rank Adaptation (hu2022lora, LoRA) is used for parameter-efficient adaptation. We use 10% of the data held out for validation; the optimal checkpoint is selected based on validation set performance. The three models, Qwen-3-8B, InternLM-3-8B and Llama-3.1-8B, are trained for 120, 160 and 80 steps, respectively. All SFT experiments are carried out using LlamaFactory (zheng2024llamafactory) on 8 NVIDIA A800 GPUs.
@@
Evaluation Results on LexChaineval
The evaluation results for a broad range of state-of-the-art LLMs on LexChaineval are reported in Table 1. In the zero-shot setting, most models exhibit limited capability in reasoning over the multiple dimensions of tort legal analysis. The best-performing model without any task-specific prompt or fine-tuning is Qwen-3-8B, achieving an overall score of 48.79. Notably, as one of the most widely used LLMs, GPT-4o achieves only 41.17 overall, underscoring the challenge posed by the tort legal reasoning task.
We observe that the LexChain-style legal prompting approach (PromptLC) leads to consistent performance improvements across all models. The largest gain is observed for DeepSeek-R1, achieving a 34.1% increase over its zero-shot baseline and yielding the highest overall performance among all tested approaches. These results highlight the effectiveness of explicit, structured legal prompting in eliciting enhanced legal reasoning abilities from current models.
We further examine the impact of fine-tuning with reasoning-enhanced data, comparing models trained on syllogistic reasoning data (SFTSyll), LexChain-style reasoning data (SFTLC) and LexChain-style preference data (DPOLC). We find that SFTSyll yields reduced performance compared to zero-shot baselines for all models. This suggests that generic syllogistic reasoning is insufficient to capture the domain-specific structure and nuances in tort reasoning. In contrast, both SFTLC and DPOLC exhibit consistent performance improvements across models, with significant gains [^3] Significance is tested through paired t-test ($\alpha$=0.01). observed for Qwen-3-8B and Llama-3.1-8B.
@@
Figure 3: Evaluation results of GPT-4o and DeepSeek-R1 on the four most common types of tort liability, including liability for motor vehicle traffic accident (TA), liability for injury to providers of labor services (LS), liability for product (P) and liability for medical damage (MD).
@@
To evaluate the generalizability of our proposed reasoning-incorporated approaches to other tasks, we conduct experiments on two related legal AI tasks: Legal Named-Entity Recognition and Criminal Damages Calculation (fei2024lawbench). Although both tasks are situated within the context of criminal cases, they share certain similarities with the legal element identification and the compensation calculation components of our LexChain framework. We compare the performance of reasoning-enhanced models trained using SFT and DPO to their zero-shot counterparts. As shown in Table 2, both the SFT-based and DPO-based approaches consistently achieve comparable or better performance on these tasks, suggesting that explicitly integrating legal reasoning chains into model training can enhance model performance across a broader spectrum of legal tasks.
@@
To analyze model performance across different tort categories, we classify test instances in LexChaineval according to their underlying dispute types. We focus on the four most prevalent types of tort liability in real-world legal practice (cjo2013): liability for motor vehicle traffic accident, liability for injury to providers of labor services, liability for product and liability for medical damage, and examine model performance within each category.
@@ Conclusion
In this work, we propose LexChain, a structured reasoning framework that explicitly models legal reasoning for Chinese tort-related cases. We introduce a benchmark to evaluate LLMs’ ability to identify key reasoning steps in tort case analysis, and find that existing models exhibit notable limitations on this task. By incorporating LexChain-style reasoning through legal prompting, supervised fine-tuning and direct preference optimization, we observe consistent performance improvements on tort reasoning. These reasoning-enhanced models also generalize well to other legal AI tasks. These results highlight the importance of explicitly modeling legal reasoning structures to enhance LLMs’ ability for authentic, domain-sensitive legal analysis.
Acknowledgments
We thank the anonymous reviewers for their insightful feedback and suggestions. We are deeply grateful to the legal practitioners we consulted in the early stage of this research. Their generous insights and practical perspectives helped shape the legal reasoning framework proposed in this paper. This work is also supported by the AI9Stars community.